OpenStudy (anonymous):

a point moves on the parabola y^2=8 in such a way that the rate of change of the ordinate is always 5 units/second. how fast is the abscissa changing when the ordinate is 4?

OpenStudy (anonymous):

Thanks you for showing up D:

OpenStudy (badhi):

shouldnt there be an x term in the parabola?

OpenStudy (anonymous):

Oh yea

OpenStudy (badhi):

so in general what you should do is differentiate the equation wrt to time. For example if the equation is \(x = f(y)\) \[\frac {dx}{dt} = \frac{df(y)}{dt}\] which further can be represented as, \[\frac{dx}{dt} = \frac{df(y)}{dy}\frac{dy}{dt}\] so when \(y=4\) find the \(\frac{dx}{dt}\) \[\left[\frac{dx}{dt} \right]_{y=4} = \left[\frac{df(y)}{dy}\right]_{y=4} \left[\frac{dy}{dt}\right]_{y=4} \]

OpenStudy (anonymous):

okay, so I have to find that?

OpenStudy (badhi):

yeah find the values of the expressions in the r.h.s

OpenStudy (anonymous):

kk one second

OpenStudy (anonymous):

\[ \frac{ ds }{ dt }=5?\]

OpenStudy (badhi):

but according to your given problem, the question itself contradicts. since it is stated that \(y^2=8\implies y = \pm \sqrt 8\) which is a constant and later it talks about a rate of change in the value of y wrt time.. Can you state the equation correctly

OpenStudy (anonymous):

I have no clue about this :/

OpenStudy (badhi):

are you sure that the question given to is like that?

OpenStudy (anonymous):

I think i may be able to figure it out, but thank you soooo much man! :D

OpenStudy (anonymous):

medal for you ;)

OpenStudy (badhi):

u are welcome ;)

OpenStudy (anonymous):

=D